OpenStudy (anonymous):

how do you join the fraction (sinx/cosx)^2(1+(cosx/sinx)^2)

OpenStudy (mertsj):

\[(\frac{\sin ^2x}{\cos ^2x})(1+\frac{\cos^2 x}{\sin^2 x})\]

OpenStudy (mertsj):

Use the distributive property

OpenStudy (anonymous):

thank you